Use the following to answer questions 32-36 Tony Stark has decided his job is too stressful and he needs to take some time off from saving the world. After reading an article in the Wall Street Journal, Mr. Stark has decided that running a hot dog stand would be a perfect fit. Mr Stark has designed and manufactured his own cart, using only scrap parts he had sitting around his workshop. He estimates he could sell the cart for S5.000. Necessary licenses and fees will cost $2,000. He plans to pay a sign spinner S15/hour to stand on the street corner. Hot dog buns cost 10 cents each and hot dogs will cost him 60 cents each. Tony Stark plans to sell his Ironman Dogs for $4 each. Use this information to answer the following questions 32. Which of the Mr Starks would be considered variable? A) The licenses and fees B) The sign spinner C) Hot dogs and hot dog buns D) Only the sign spinner and the hot dogs and buns E) All of the above are variable costs 33. Assume he has not purchased hot dogs, buns or hired a sign spinner yet. At this point, how much are Mr. Starks fixed costs? A) S0 B) $2,000 C) $5,000 D) $7,000 34. Tony Stark has decided his hot body is enough to drive business and is not going to employ a sign spinner or do any other advertising. How many hot-dogs will he need to sell to break-even? (Round up to the nearest whole number) A) 500 B) 606 C) 1,750 D) 2122
Answer #1
Question Answer Explanation
32 D Variable costs are variable in nature. License fees is fixed cost.
33 B Only license fees is the cost.
34 B =2000/(4-0.6-0.1)=606
35 C =1000*(0.6+0.1)+2000=2,700
Please raise another question for remaining part as per Chegg policy.